OpenStudy (anonymous):

A mother puts $10,000 into a special bank account on the date of birth of her first born. The account earns interest at a rate of 4%, and compounds 4 times per year. When the child turns 18, he receives the lump sum. How much will the child receive? A. $10,407.65 B. $11,961.47 C. $20,470.99 D. $168,422.62

OpenStudy (anonymous):

@sreekar369 @suckerofmath @sikinder

OpenStudy (anonymous):

The answer is C.

OpenStudy (tkhunny):

Single Deposit Single Interest Rate i = 0.04 m = 4 j = i/m = 0.01 n = 18 n*m = 18*4 = 72 10000*(1 + 0.01)^{72} = ?